OpenStudy (anonymous):

Find the integral of sin(x^2) from 0 to 1 using McLauren Series.

OpenStudy (anonymous):

i am not sure, but it guess you are supposed to write the series for \(\sin(x^2)\), then integrate term by term for some number of terms

OpenStudy (anonymous):

\[\sin(x)=x-\frac{x^3}{3!}+\frac{x^5}{5!}-...\] replace \(x\) by \(x^2\)